You get a new credit card from your bank. The document that comes with the card informs you that the interest rate on that card is 6.1% APR. What is the effective annual rate you'll actually be paying?
Answer #1

Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.0610/1)^1 -1

        = 1.0610 -1

        = 0.0610*100

        = 6.10%.

Therefore, the effective annual rate to be paid on the anew credit card is 6.10%.

In case of any query, kindly comment on the solution.